Determine ΔH for the following reaction.

C4H10 (g) + oxygen (g) carbon dioxide (g) + water (g
1
Expert's answer
2020-10-22T14:18:36-0400

Solution:

The balanced chemical reaction:

C4H10(g) + 6.5O2(g) = 4CO2(g) + 5H2O(g)


Table of standard enthalpy of formation:

(https://www.thoughtco.com/common-compound-heat-of-formation-table-609253)

ΔHf(C4H10) = -124.7 kJ/mol

ΔHf(CO2) = -393.5 kJ/mol

ΔHf(H2O) = -241.8 kJ/mol

ΔHf(O2) = 0 kJ/mol


Hence,

ΔHr = 5 × ΔHf(H2O) + 4 × ΔHf(CO2) - ΔHf(C4H10) - 6.5 × ΔHf(O2)

ΔHr = (5 × -241.8) + (4 × -393.5) - (-124.7) - (6.5 × 0) = -2658.3 kJ/mol

ΔHr = -2658.3 kJ/mol


Answer: ΔH for the butane (C4H10) combustion reaction is -2658.3 kJ/mol.